# ResQ: Mixed-Precision Quantization of Large Language Models with Low-Rank Residuals

### v1

@@ Abstract
Post-training quantization (PTQ) of large language models (LLMs) holds the promise in reducing the prohibitive computational cost at inference time. Quantization of all weight, activation and key-value (KV) cache tensors to 4-bit without significantly degrading generalizability is challenging, due to the high quantization error caused by extreme outliers in activations. To tackle this problem, we propose ResQ, a PTQ method that pushes further the state-of-the-art. By means of principal component analysis (PCA), it identifies a low-rank subspace (in practice $\nicefrac{{1}}{{8}}$ of the hidden dimension) in which activation variances are highest, and keep the coefficients within this subspace in high precision, e.g. 8-bit, while quantizing the rest to 4-bit. Within each subspace, invariant random rotation is applied to further suppress outliers. We show that this is a provably optimal mixed precision quantization scheme that minimizes error. With the Llama families of models, we demonstrate that ResQ outperforms recent uniform and mixed precision PTQ methods on a variety of benchmarks, achieving up to 33% lower perplexity on Wikitext than the next best method SpinQuant, and a 2.4$\times$ speedup over 16-bit baseline. Code is available at github.com/utkarsh-dmx/project-resq.
ResQ: Mixed-Precision Quantization of Large Language Models
with Low-Rank Residuals
Utkarsh Saxena1 Sayeh Sharify2, Kaushik Roy1, Xin Wang2 1Department of Electrical and Computer Engineering, Purdue University 2d-Matrix, Santa Clara, USA 1{saxenau, kaushik}@purdue.edu, 2{sayehs, xwang}@d-matrix.ai
@@ 1 Introduction
Quantization algorithms are powerful and principled approaches to address the immense computational demands of LLMs at both stages of inference. Quantization of weights reduces parameter storage, lowers memory usage of KV cahe during generation, whereas activation quantization decreases the complexity of floating-point operation. However, effective low-precision quantization is difficult due to large outliers in activations, which can be ~20×20$\times$20 × larger than other values Dettmers et al. (2022). While post-training methods like KIVI Liu et al. (2024b) and KVQuant Hooper et al. (2024) achieve 2-bit KV cache quantization, and techniques like GPTQ Frantar et al. (2022) and AWQ Lin et al. (2024c) optimize very low-precision weights, quantizing activations below 8-bit precision remains an open challenge.
@@
Recent LLM activation quantization methods feature two useful strategies: differential treatment of outliers retain outlier channels in high precision, leading to mixed-precision quantization (e.g., Dettmers et al. 2022; Zhao et al. 2024; Ashkboos et al. 2023; Figure 1a), whereas invariant random rotation suppress outliers, leading to less difficult uniform low-precision quantization (e.g., Ashkboos et al. 2024b; Liu et al. 2024a; Figure 1b). Both reduce quantization error and improve signal-to-quantization-noise ratio (Figure 1d,e) locally; yet a notable model performance gap persists from the 16-bit baseline. For example, SpinQuant Liu et al. (2024a) at 4-bit, applied to Llama-3-8b Meta (2024a), exhibits ~20% higher perplexity than 16-bit baseline, even after extensive optimization.
@@
To mend this gap, we introduce ResQ, a novel quantization method that combines the strengths of both aforementioned strategies and thereby improve model efficiency with aggressive 4-bit quantization of activation, weight, and KV cache. Specifically, by means of principal component analysis (PCA), we first identify a low-rank subspace that captures highest variances in activation, and mark the coefficients along this subspace for high-precision (8-bit) and the complement subspace for low-precision (4-bit) quantization. We prove that so doing minimizes quantization error. Moreover, ResQ employs invariant random rotations within each subspace before quantization to further suppress outliers (Figure 1c,d,e). Similar to SpinQuant, most projection matrices can be fused into adjacent weights, leading to minimal runtime computational overhead (Section 4.3). Furthermore, ResQ can be applied to KV cache quantization as well, and can be combined with GPTQ Frantar et al. (2022), resulting in even better generalizing LLMs.
@@
In practice, one could naïvely combine outlier-based and rotation-based quantization methods. For example, high-precision outliers could be detected using the $\ell_{\infty}$ norm similar to QUIK Ashkboos et al. (2023), and random rotations could be applied within both high- and low-precision quantization groups, as in QuaRot Ashkboos et al. (2024b). However, such approaches fare less well than ResQ (Figure 1d and 1e) in practice, in support of ResQ’s provably optimal treatment of outlier quantization. When quantizing weight, activation and KV cache to 4-bit with only $\nicefrac{{1}}{{8}}$ channels in 8-bit, ResQ achieves 2-33% lower perplexity on Wikitext and 1-5.5% zero-shot accuracy improvements over SpinQuant Liu et al. (2024a), the best competing method in practice. Unlike SpinQuant, ResQ does not require gradient-based optimization, making it an easier and faster PTQ solution. Additionally, the rank $r$ of the high-precision subspace can be tuned, giving rise to Pareto-optimal solutions as a tradeoff between computational efficiency and accuracy.
We claim the following contributions.
@@
    We develop hardware kernels and achieve runtime speedup on NVIDIA A100 GPU with our quantized models.
@@ 2.1 Quantization of LLMs
Quantization reduces model size and accelerates inference by lowering neural network bit precision (Choi et al., 2018; Hubara et al., 2021; Yao et al., 2022; Park et al., 2022; Gholami et al., 2022; Xi et al., 2023). It is broadly categorized into two categories: uniform precision quantization (UPQ) and mixed precision quantization (MPQ).
Uniform precision quantization (UPQ) applies the same bit-width across all layers, simplifying implementation but neglecting layer-specific sensitivity to quantization.
Weight-only UPQ methods reduce storage by compressing weights, using techniques like Hessian-guided rounding (GPTQ, Frantar et al. 2022), adaptive rounding (QuIP, Chee et al. 2024), channel-wise scaling (AWQ, Lin et al. 2024c), and multi-codebook quantization (AQLM, Egiazarian et al. 2024). However, these methods struggle with batch processing due to significant activation memory overhead. Weight-activation UPQ compresses both weights and activations to address this. Methods such as SmoothQuant (Xiao et al., 2023) and OmniQuant (Shao et al., 2023) scale activations and weights to handle outliers, while RPTQ (Yuan et al., 2023a), QLLM (Liu et al., 2023a), and QServe (Lin et al., 2024d) employ channel-level strategies like clustering and reordering. Rotation-based methods such as QuaRot (Ashkboos et al., 2024b), SpinQuant (Liu et al., 2024a) and DuQuant Lin et al. (2024b) further enhance robustness in low-precision scenarios. KV cache UPQ reduces memory for large batches or long contexts. FlexGen (Sheng et al., 2023) employs 4-bit quantization and memory offloading, while KIVI (Liu et al., 2024b) uses asymmetric 2-bit quantization for compression, enabling efficient inference.
@@
optimizes bit-widths by adapting to the sensitivity of weights and activations, achieving better accuracy than UPQ at similar compression rates. Our proposed method, ResQ, follows the MPQ approach.
Weight-only MPQ has advanced efficiency for memory-bound applications with minimal activation demands. Methods like OWQ (Lee et al., 2024) and SpQR (Dettmers et al., 2023) mitigate activation outliers’ impact by retaining critical features in full precision, while SqueezeLLM (Kim et al., 2023) employs Dense-and-Sparse decomposition to efficiently store sensitive weights. Weight-activation MPQ enhances efficiency by addressing activation outliers (e.g. Guan et al. 2024; Zeng et al. 2024). Methods like LLM.int8()(Dettmers et al., 2022) and QUIK(Ashkboos et al., 2023) preserve critical activations with mixed or low-precision decompositions, while Atom (Zhao et al., 2024) and SliM-LLM (Huang et al., 2024) optimize quantization through channel reordering and salience-driven bit allocation. KV cache MPQ reduces memory usage while preserving precision for critical tokens using techniques like non-uniform quantization, importance-aware precision, and salient token compression (Hooper et al., 2024; Yang et al., 2024; Dong et al., 2024; He et al., 2024). Alternatively, GEAR quantizes all tokens’ KV cache and maintains quantization error in a low-rank form (Kang et al., 2024).
@@ 2.2 Low-rank Decomposition
Low-rank decomposition techniques have been widely used in model compression, reducing dimensionality while maintaining performance. For instance, SliceGPT Ashkboos et al. (2024a) projects weight matrices onto principal components for sparsification, while ESPACE Sakr and Khailany (2024) reduces activation dimensionality via pre-calibrated projections, achieving inference-time efficiency. Similarly, ASVD Yuan et al. (2023b) introduces an activation-aware decomposition method that incorporates activation distributions into weight decomposition.
Additionally, low-rank decomposition can be applied to reduce KV cache size. For example, Eigen Attention Saxena et al. (2024) and ASVD Yuan et al. (2023b) employ low-rank approximations to reduce memory usage in KV caches during attention operations. PALU Chang et al. (2024) introduces learnable projections to adaptively compress KV caches based on the compression budget. Finally, Matryoshka KV Cache refines this with hierarchical orthogonal projections and knowledge distillation Lin et al. (2024a).
@@ 4.1 Quantization scheme
Given input activation ${\bm{X}}\in{\mathbb{R}}^{n\times d}$ and weight ${\bm{W}}\in\mathbb{R}^{d\times d}$, they are first projected onto an orthogonal basis defined by the vectors ${\bm{U}}\in{\mathbb{R}}^{d\times d}$. The coefficients of the projections along this basis are then subject to quantization. We seek to quantize some coefficients along certain bases at high precision while those remaining at low precision. Within $\mathbb{R}^{d}$, denote bases of a low-rank space of high-precision components by ${\bm{U}}_{h}\in{\mathbb{R}}^{d\times r}$ and those of its complementary subspace of low-precision components by ${\bm{U}}_{l}\in{\mathbb{R}}^{d\times(d-r)}$. The rank $r$ controls the amount of components in high precision (in practice we typically choose $r=\nicefrac{{d}}{{8}}$). We have ${\bm{U}}_{h}{\bm{U}}^{\top}_{h}+{\bm{U}}_{l}{\bm{U}}^{\top}_{l}={\bm{U}}{\bm{U}}^{\top}={\bm{I}}$ because ${\bm{U}}$ is orthogonal.
The quantized activation ${{\bm{X}}_{q}}$ is thusly
${{\bm{X}}_{q}}=Q_{L}({{\bm{X}}{\bm{U}}_{l}}){{\bm{U}}^{\top}_{l}}+Q_{H}({{\bm{X}}{\bm{U}}_{h}}){{\bm{U}}^{\top}_{h}}$ (2)
Similarly, quantized weights ${\bm{W}}_{q}$ is obtained by projecting the inputs space of weights and quantizing the coefficients,
${{\bm{W}}_{q}}={{\bm{U}}_{l}}Q_{L}({{\bm{U}}^{\top}_{l}{\bm{W}}})+{{\bm{U}}_{h}}Q_{H}({{\bm{U}}^{\top}_{h}{\bm{W}}}).$ (3)
@@
4.2 Projections and optimality thereof
The orthogonal basis vectors ${\bm{U}}$ should intuitively have two properties: (1) the low-rank space for high-precision quantization should capture the more important components, and (2) quantization error in both high- and low-precision groups should be minimized. We construct ${\bm{U}}$ as a combination of two rotation matrices serving both objectives respectively. We write ${{\bm{U}}_{i}={\bm{P}}_{i}{\bm{R}}_{i}},i\in{\{h,l\}}$. Therefore,
where, ${{\bm{P}}_{l},{\bm{R}}_{l}}\in{{\mathbb{R}}}^{d\times(d-r)},{{\bm{P}}_{h},{\bm{R}}_{h}}\in{{\mathbb{R}}}^{d\times r}$. Inspired by prior work Ashkboos et al. (2024b); Chee et al. (2024), we make ${{\bm{R}}_{l},{\bm{R}}_{h}}$ random orthogonal matrices because random rotation reduces outliers, making the rotated matrices easier to quantize. Furthermore, projecting data using a random orthogonal matrix increases Gaussianity of the activations and weights within high- and low-precision groups, due to Lemma 4.1, conducive to the quantizations applied to these groups.
@@ 4.2 Projections and optimality thereof
By Central Limit Theorem, the distribution after multiplication with random orthogonal matrix follows an approximately bell-shaped Gaussian distribution Chee et al. (2024).
@@
Proof of Theorem 7 can be found Appendix A.1. Theorem 7 bounds the quantization error in Equation 6. To lower the upper bound of quantization error, it is required to maximize $\lVert{\bm{X}}{\bm{P}}_{h}\rVert_{F}$ which happens when ${\bm{P}}_{h}$ comprises of eigenvectors of the covariance matrix ${\bm{X}}{\bm{X}}^{\top}$ with its largest eigenvalues. Therefore the low-rank subspace for high-precision quantization can be obtained by means of PCA, while the subspace for low-precision quantization can be obtained using ${{\bm{U}}_{h}{\bm{U}}^{\top}_{h}+{\bm{U}}_{l}{\bm{U}}^{\top}_{l}={\bm{P}}_{h}{\bm{P}}^{\top}_{h}+{\bm{P}}_{l}{\bm{P}}^{\top}_{l}={\bm{I}}}$ (because ${{\bm{R}}_{i}}$ is orthogonal). If we construct ${\bm{P}}$ by taking eigenvectors of ${\bm{X}}{\bm{X}}^{\top}$ arranged in increasing order of eigenvalues, the last $r$ columns of such a ${\bm{P}}$ would correspond to ${\bm{P}}_{h}$ and the first $d-r$ columns would correspond to ${\bm{P}}_{l}$. Appendix A.2 shows the distribution of activations after projecting them to the space ${{\bm{U}}}$.

### v2

@@ Abstract
Post-training quantization (PTQ) of large language models (LLMs) holds the promise in reducing the prohibitive computational cost at inference time. Quantization of all weight, activation and key-value (KV) cache tensors to 4-bit without significantly degrading generalizability is challenging, due to the high quantization error caused by extreme outliers in activations. To tackle this problem, we propose ResQ, a PTQ method that pushes further the state-of-the-art. By means of principal component analysis (PCA), it identifies a low-rank subspace (in practice $\nicefrac{{1}}{{8}}$ of the hidden dimension) in which activation variances are highest, and keep the coefficients within this subspace in high precision, e.g. 8-bit, while quantizing the rest to 4-bit. Within each subspace, invariant random rotation is applied to further suppress outliers. We show that this is a provably optimal mixed precision quantization scheme that minimizes error. With the Llama and Qwen2.5 families of models, we demonstrate that ResQ outperforms recent uniform and mixed precision PTQ methods on a variety of benchmarks, achieving up to 33% lower perplexity on Wikitext than the next best method SpinQuant, and upto 3$\times$ speedup over 16-bit baseline. Code repository available here.  [^1] https://github.com/utkarsh-dmx/project-resq
Machine Learning, ICML
@@ 1 Introduction
Quantization algorithms are powerful and principled approaches to address the immense computational demands of LLMs at both stages of inference. Quantization of weights reduces parameter storage, KV cache quantization mlowers memory usage of KV cache during generation, whereas activation quantization decreases the complexity of floating-point operation. However, effective low-precision quantization is difficult due to large outliers in activations, which can be 20$\sim$ 20$\times$ 20 × larger than other values (Dettmers et al., 2022). While post-training methods like KIVI (Liu et al., 2024a) and KVQuant (Hooper et al., 2024) achieve 2-bit KV cache quantization, and techniques like GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2024c) optimize very low-precision weights, quantizing activations below 8-bit precision remains an open challenge.
@@
Recent LLM activation quantization methods feature two useful strategies: differential treatment of outliers retain outlier channels in high precision, leading to mixed-precision quantization (e.g., Dettmers et al. 2022; Zhao et al. 2024; Ashkboos et al. 2024b; Figure 1a), whereas invariant random rotation suppress outliers, leading to less difficult uniform low-precision quantization (e.g., Ashkboos et al. 2024c; Liu et al. 2024b; Figure 1b). Both reduce quantization error and improve signal-to-quantization-noise ratio (Figure 1d,e) locally; yet a notable model performance gap persists from the 16-bit baseline. For example, SpinQuant (Liu et al., 2024b) at 4-bit, applied to Meta-Llama-3-8B (Meta, 2024b), exhibits $\sim 20\%$ higher perplexity than the 16-bit floating point baseline, even after nontrivial optimization.
@@
To mend this gap, we introduce ResQ, a novel PTQ method that combines the strengths of both aforementioned strategies and thereby improve model efficiency with aggressive 4-bit quantization of activation, weight, and KV cache. Specifically, by means of principal component analysis (PCA), we first identify a low-rank subspace that captures highest variances in activation, and mark the coefficients along this subspace for high-precision (8-bit) and the complement subspace for low-precision (4-bit) quantization. Then, ResQ employs invariant random rotations within each subspace before quantization to further suppress outliers (Figure 1c,d,e). We prove that the above treatment minimizes quantization error. Similar to SpinQuant, most projection matrices can be fused into adjacent weights, leading to minimal runtime computational overhead (Section 4.3). Furthermore, ResQ can be applied to KV cache quantization as well, and can be combined with GPTQ (Frantar et al., 2022), resulting in even better generalizing LLMs.
@@
Outlier-based and rotation-based quantization methods can be combined. For example, high-precision outliers could be detected by $\ell_{\infty}$-norm similar to QUIK (Ashkboos et al., 2024b), and random rotations applied within both high- and low-precision quantization groups, as in QuaRot (Ashkboos et al., 2024c). These methods fare less well than ResQ (Figure 1d,e) in practice, in support of ResQ’s provably optimal treatment of outlier quantization. When quantizing weight, activation and KV cache to 4-bit with only $\nicefrac{{1}}{{8}}$ channels in 8-bit, ResQ achieves 4-33% lower perplexity on Wikitext and 0.1-5.4% 0-shot accuracy improvements over SpinQuant (Liu et al., 2024b), the best in practice so far. Unlike SpinQuant, ResQ does not require gradient-based optimization, making it a less demanding and faster PTQ solution. Furthermore, tuning the rank $r$ of ResQ gives rise to Pareto-optimal solutions as a tradeoff between efficiency and accuracy. We claim the following contributions.
@@
    We develop CUDA kernels and achieve runtime speedup on NVIDIA GPUs with our quantized models.
@@ 2.1 Quantization of LLMs
Quantization reduces model size and accelerates inference by lowering neural network bit precision (Choi et al., 2018; Hubara et al., 2021; Yao et al., 2022; Park et al., 2022; Gholami et al., 2022; Xi et al., 2023). It is broadly categorized into two categories: uniform precision quantization (UPQ) and mixed precision quantization (MPQ). Uniform precision quantization (UPQ) applies the same bit-width across all layers, simplifying implementation but neglecting layer-specific sensitivity to quantization. Weight-only UPQ methods reduce storage by compressing weights, using techniques like Hessian-guided rounding (GPTQ, Frantar et al. 2022), adaptive rounding (QuIP, Chee et al. 2024), channel-wise scaling (AWQ, Lin et al. 2024c), and multi-codebook quantization (AQLM, Egiazarian et al. 2024). However, these methods struggle with batch processing due to significant activation memory overhead. Weight-activation UPQ compresses both weights and activations to address this. Methods such as SmoothQuant (Xiao et al., 2023) and OmniQuant (Shao et al., 2023) scale activations and weights to handle outliers, while RPTQ (Yuan et al., 2023a), QLLM (Liu et al., 2023a), and QServe (Lin et al., 2024d) employ channel-level strategies like clustering and reordering. Rotation-based methods such as QuaRot (Ashkboos et al., 2024c), SpinQuant (Liu et al., 2024b) and DuQuant (Lin et al., 2024b) further enhance robustness in low-precision scenarios. KV cache UPQ reduces memory for large batches or long contexts. FlexGen (Sheng et al., 2023) employs 4-bit quantization and memory offloading, while KIVI (Liu et al., 2024a) uses asymmetric 2-bit quantization for compression, enabling efficient inference.
@@
optimizes bit-widths by adapting to the sensitivity of weights and activations, achieving better accuracy than UPQ at similar compression rates. Our proposed method, ResQ, follows the MPQ approach. Weight-only MPQ has advanced efficiency for memory-bound applications with minimal activation demands. Methods like OWQ (Lee et al., 2024) and SpQR (Dettmers et al., 2023) mitigate activation outliers’ impact by retaining critical features in full precision, while SqueezeLLM (Kim et al., 2023) employs Dense-and-Sparse decomposition to efficiently store sensitive weights. Weight-activation MPQ enhances efficiency by addressing activation outliers (e.g. Guan et al. 2024; Zeng et al. 2024). Methods like LLM.int8() (Dettmers et al., 2022) and QUIK(Ashkboos et al., 2024b) preserve critical activations with mixed or low-precision decompositions, while Atom (Zhao et al., 2024) and SliM-LLM (Huang et al., 2024) optimize quantization through channel reordering and salience-driven bit allocation. KV cache MPQ reduces memory usage while preserving precision for critical tokens using techniques like non-uniform quantization, importance-aware precision, and salient token compression (Hooper et al., 2024; Yang et al., 2024b; Dong et al., 2024; He et al., 2024). Alternatively, GEAR quantizes all tokens’ KV cache and maintains low-rank quantization error (Kang et al., 2024).
@@ 2.2 Low-rank decomposition
Low-rank decomposition techniques have been widely used in model compression, reducing dimensionality while maintaining performance. For instance, SliceGPT (Ashkboos et al., 2024a) projects weight matrices onto principal components for sparsification, while ESPACE (Sakr & Khailany, 2024) reduces activation dimensionality via pre-calibrated projections, achieving inference-time efficiency. Similarly, ASVD (Yuan et al., 2023b) introduces an activation-aware decomposition method that incorporates activation distributions into weight decomposition. Additionally, low-rank decomposition can be applied to reduce KV cache size. For example, Eigen Attention (Saxena et al., 2024) and ASVD (Yuan et al., 2023b) employ low-rank approximations to reduce memory usage in KV caches during attention operations. PALU (Chang et al., 2024) introduces learnable projections to adaptively compress KV caches based on the compression budget. Finally, Matryoshka KV Cache refines this with hierarchical orthogonal projections and knowledge distillation (Lin et al., 2024a).
@@ 4.1 Quantization scheme
Given input activation ${\bm{X}}\in{\mathbb{R}}^{n\times d}$ and weight ${\bm{W}}\in\mathbb{R}^{d\times d}$, they are first projected onto an orthogonal basis defined by the vectors ${\bm{U}}\in{\mathbb{R}}^{d\times d}$. The coefficients of the projections along this basis are then subject to quantization. We seek to quantize some coefficients along certain bases at high precision while those remaining at low precision. Within $\mathbb{R}^{d}$, denote bases of a low-rank space of high-precision components by ${\bm{U}}_{h}\in{\mathbb{R}}^{d\times r}$ and those of its complementary subspace of low-precision components by ${\bm{U}}_{l}\in{\mathbb{R}}^{d\times(d-r)}$. The rank $r$ controls the amount of components in high precision (in practice we typically choose $r=\nicefrac{{d}}{{8}}$). We have ${\bm{U}}_{h}{\bm{U}}^{\top}_{h}+{\bm{U}}_{l}{\bm{U}}^{\top}_{l}={\bm{U}}{\bm{U}}^{\top}={\bm{I}}$ because ${\bm{U}}$ is orthogonal. The quantized activation ${{\bm{X}}_{q}}$ is thusly
${{\bm{X}}_{q}}=Q({{\bm{X}}{\bm{U}}})=Q_{L}({{\bm{X}}{\bm{U}}_{l}})+Q_{H}({{\bm{X}}{\bm{U}}_{h}})$ (2)
Similarly, quantized weights ${\bm{W}}_{q}$ is obtained by projecting the inputs space of weights by ${\bm{U}}^{\top}$ and quantizing the coefficients,
${{\bm{W}}_{q}}=Q({{\bm{U}}^{\top}{\bm{W}}})=Q_{L}({{\bm{U}}^{\top}_{l}{\bm{W}}})+Q_{H}({{\bm{U}}^{\top}_{h}{\bm{W}}}).$ (3)
@@
4.2 Projections and optimality thereof
Intuitively, the orthogonal basis vectors ${\bm{U}}$ should have two properties: (1) the low-rank space for high-precision quantization should capture the more important components, and (2) quantization error in both high- and low-precision groups should be minimized. We construct ${\bm{U}}$ as a combination of two rotation matrices serving both objectives respectively. We write ${{\bm{U}}_{i}={\bm{P}}_{i}{\bm{R}}_{i}},i\in{\{h,l\}}$. Therefore,
where, ${{\bm{P}}_{l},{\bm{R}}_{l}}\in{{\mathbb{R}}}^{d\times(d-r)},{{\bm{P}}_{h},{\bm{R}}_{h}}\in{{\mathbb{R}}}^{d\times r}$. Inspired by prior work (Ashkboos et al., 2024c; Chee et al., 2024), we make ${{\bm{R}}_{l},{\bm{R}}_{h}}$ random orthogonal matrices because random rotation reduces outliers, making the rotated matrices easier to quantize. Furthermore, projection with a random orthogonal matrix increases Gaussianity of activations and weights within high- and low-precision groups, due to Lemma 4.1, conducive to the quantizations applied to these groups.
Figure 3: Activation distribution of the baseline and applying the projection matrices.
@@ 4.2 Projections and optimality thereof
By Central Limit Theorem, the distribution after multiplication with random orthogonal matrix is approximately Gaussian (Tseng et al., 2024).
@@
Full proof of Theorem 7 is in Appendix A. Theorem 7 bounds the quantization error in Equation 6 from above. To lower this upper bound of quantization error is thusly to maximize $\lVert{\bm{X}}{\bm{P}}_{h}\rVert_{F}$ which happens when ${\bm{P}}_{h}$ comprises of eigenvectors of the covariance matrix ${\bm{X}}{\bm{X}}^{\top}$ with its largest eigenvalues. Therefore, the low-rank subspace for high-precision quantization can be obtained by means of PCA, while the subspace for low-precision quantization can be obtained using ${{\bm{U}}_{h}{\bm{U}}^{\top}_{h}+{\bm{U}}_{l}{\bm{U}}^{\top}_{l}={\bm{P}}_{h}{\bm{P}}^{\top}_{h}+{\bm{P}}_{l}{\bm{P}}^{\top}_{l}={\bm{I}}}$ (because ${{\bm{R}}_{i}}$ is orthogonal). If we construct ${\bm{P}}$ by taking eigenvectors of ${\bm{X}}{\bm{X}}^{\top}$ arranged in increasing order of eigenvalues, the last $r$ columns of such a ${\bm{P}}$ would correspond to ${\bm{P}}_{h}$ and the first $d-r$ columns would correspond to ${\bm{P}}_{l}$. The distribution of activation after applying different projection matrices is shown in Figure 3. Projection of activation along ${\bm{P}}$ sorts the activation coefficients in increasing order of variance due to increasing eigenvalues of bases vectors. Consequently, the later $r$ channels of the projected activations with higher variance are kept in higher precision. Projection along ${\bm{U}}={\bm{P}}{\bm{R}}$ smoothes the activations along low precision and high precision groups further reducing quantization error (Figure 3) and improving quantization SNR (Figure 1(d,e)).